OpenStudy (anonymous):

Without a calculator, determine the exact value of each expression. sin (π/2) cos (π/3) Could someone please explain to me how this problem is solved. I know I did this in Pre-Cal, but I honestly don't remember a thing. Thanks.

jimthompson5910 (jim_thompson5910):

do you remember the unit circle?

OpenStudy (littlebird):

Sin(pi/2)=Sin(90) I think sin was the y values y value at sin 90= 1 Think of the unit circle

OpenStudy (anonymous):

what @jim_thompson5910 said there is a nice unit circle on the last page of the attached cheat sheet find \(\frac{\pi}{3}\)and look at the corresponding point the first coordinate is the cosine and the second coordinate is the sine

OpenStudy (anonymous):

now what do I do with the unit circle? btw, I need to show all work and + I need the work to understand it myself. There's about 14questions like these two, but I figured if I could understand these two I could do the rest on my own.

jimthompson5910 (jim_thompson5910):

do you see on either unit circle given where it has pi/3 as the angle?

OpenStudy (anonymous):

i can see why you need to understand this but there is no work to show it is not a computation like \(7^2\) or something look at the point on the unit circle corresponding to the angle \(\frac{\pi}{3}\)

OpenStudy (anonymous):

you should see the point \((\frac{1}{2},\frac{\sqrt3}{2})\)

OpenStudy (anonymous):

the first coordinate is cosine, and the second coordinate is sine i.e \[\cos(\frac{\pi}{3})=\frac{1}{2}\\ \sin(\frac{\pi}{3})=\frac{\sqrt3}{2}\]

OpenStudy (anonymous):

so there is no actual work to "show" all the work is done with your eyeballs, locating the point you need then you write the answer

OpenStudy (littlebird):

When I took calculus my teacher tested us without providing us with a unit circle. This website might help you memorize it. http://www.mathsisfun.com/geometry/unit-circle.html

OpenStudy (anonymous):

Would it be acceptable if for the work I did- Step 1 Find the reference angle, θ' and sin θ' θ' = 180 - 90 = 90 The function value for the reference angle is sin 90 = 1

jimthompson5910 (jim_thompson5910):

that will work, but you can also just state that sin(90) = 1 so sin(pi/2) = 1 as satellite73 is saying, there's not much work to show since you're looking up a table or a diagram (or just stating what you have memorized)

jimthompson5910 (jim_thompson5910):

also, here's a great page for reference angles http://www.mathopenref.com/reference-angle.html

OpenStudy (anonymous):

I know there isn't much work, but this is the first assignment for the school year and I just want to make a good impression. Now how would I do tan (7pi/4) tan is equal y/x correct.

jimthompson5910 (jim_thompson5910):

yes because tan = sin/cos